OpenStudy (anonymous):

The area of a square garden is 128 m². How long is the diagonal?

OpenStudy (anonymous):



OpenStudy (anonymous):

SO WHTS THE ANSWER

OpenStudy (anonymous):

what do you think the answer is ? \[d = \sqrt{m^2 + m^2}=\sqrt{128 + 128}\]